a solid sphere rolls without slipping from height of 3.5m down inclined plane. calculate speed of sphere when it reaches bottom of ramp.
Answer #1

work by a done gravity change in kinetic Energy mgha Imret II wr 2 = 2/5 m² w = VIR righ = 7 62 13 14 X9883-5 6= 7 ml